# Thermodynamically consistent Cahn-Hilliard-Navier-Stokes equations using the metriplectic dynamics formalism

### 16 Feb 2024 v1

@@ 3.3 Metriplectic 4-bracket for the Cahn-Hilliard-Navier-Stokes system
$\psi(\mathbf{x},t)=(\mathbf{m}(\mathbf{x},t),\rho(\mathbf{x},t),\tilde{c}(\mathbf{x},t),\sigma^{a}(\mathbf{x},t))\,,$ (135)
@@
${\Sigma}(dF,dG)$ $=\nabla F_{{\mathbf{m}}}:\bar{\bar{\Lambda}}_{1}:\nabla G_{{\mathbf{m}}}+\nabla F_{{{\sigma}^{a}}}\cdot\bar{\Lambda}_{2}\cdot\nabla G_{{{\sigma}^{a}}}+\nabla\mathcal{L}^{a}_{\tilde{c}}(F)\cdot\bar{\Lambda}_{3}\cdot\mathcal{L}^{a}_{\tilde{c}}(F)\,,$ (137)
@@
Here the tensors $\bar{\bar{\Lambda}}_{1}$ , $\bar{\Lambda}_{2}$ and $\bar{\Lambda}_{3}$ are defined as (60) of §§ 2.3.3. Then, the 4-bracket reads
$(F,K;G,N)^{a}$ $=\int_{\Omega}\frac{1}{T}\Big{[}\left[K_{{\sigma}^{a}}\nabla F_{\mathbf{m}}-F_{{\sigma}^{a}}\nabla K_{\mathbf{m}}\right]\colon\!\bar{\bar{\Lambda}}\colon\left[N_{{\sigma}^{a}}\nabla G_{\mathbf{m}}-G_{{\sigma}^{a}}\nabla N_{\mathbf{m}}\right]$
@@
$+\big{[}K_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(F)-F_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(K)\big{]}\cdot\bar{D}\cdot\big{[}N_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(G)-G_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(N)\big{]}\,.$ (139)
Observe, with the exception of the last line this bracket, it is identical to that of (61).
@@
$=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot\big{[}p\mathbf{I}+\lambda_{f}\rho\Gamma{\bm{\xi}}\otimes\nabla c\big{]}+\frac{1}{\rho}\nabla\cdot(\bar{\bar{\Lambda}}:\nabla\mathbf{v})\,,$ (141)
$\partial_{t}\rho$ =$=\{\rho,H^{1}\}^{1}+(\rho,H^{1};S^{1},H^{1})^{1}=-\mathbf{v}\cdot\nabla\rho-\rho\,\nabla\cdot\mathbf{v}\,,$ (142)
$\partial_{t}\tilde{c}$ =$=\{\tilde{c},H^{1}\}^{1}+(\tilde{c},H^{1};S^{1},H^{1})^{1}=-\mathbf{v}\cdot\nabla\tilde{c}-\tilde{c}\,\nabla\cdot\mathbf{v}+\nabla\cdot(\bar{D}\cdot\nabla\mu^{1}_{\Gamma})\,,$ (143)
$\partial_{t}\sigma^{1}_{\mathrm{Total}}$ =$=\{\sigma^{1}_{\mathrm{Total}},H^{1}\}^{1}+(\sigma^{1}_{\mathrm{Total}},H^{1};S^{1},H^{1})^{1}=-\mathbf{v}\cdot\nabla\sigma^{1}_{\mathrm{Total}}-\sigma^{1}_{\mathrm{Total}}\,\nabla\cdot\mathbf{v}$
$\quad+\nabla\cdot\left(\frac{\bar{\kappa}}{T}\cdot\nabla T\right)+\frac{1}{T^{2}}\nabla T\cdot\bar{\kappa}\cdot\nabla T+\frac{1}{T}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla\mu^{1}_{\Gamma}\cdot\bar{D}\cdot\nabla\mu^{1}_{\Gamma}\,.$ (144)
@@
$=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot\big{[}\left(p-\lambda_{f}\Gamma^{2}/2\right)\mathbf{I}+\lambda_{f}\Gamma{\bm{\xi}}\otimes\nabla c\big{]}+\frac{1}{\rho}\nabla\cdot(\bar{\bar{\Lambda}}:\nabla\mathbf{v})\,,$ (145)
$\partial_{t}\rho$ =$=\{\rho,H^{0}\}^{0}+(\rho,H^{0};S^{0},H^{0})^{0}=-\mathbf{v}\cdot\nabla\rho-\rho\,\nabla\cdot\mathbf{v}\,,$ (146)
$\partial_{t}\tilde{c}$ =$=\{\tilde{c},H^{0}\}^{0}+(\tilde{c},H^{0};S^{0},H^{0})^{0}=-\mathbf{v}\cdot\nabla\tilde{c}-\tilde{c}\,\nabla\cdot\mathbf{v}+\nabla\cdot(\bar{D}\cdot\nabla\mu^{0}_{\Gamma})\,,$ (147)
$\partial_{t}\sigma^{0}_{\mathrm{Total}}$ =$=\{\sigma^{0}_{\mathrm{Total}},H^{0}\}^{0}+(\sigma^{0}_{\mathrm{Total}},H^{0};S^{0},H^{0})^{0}=-\mathbf{v}\cdot\nabla\sigma^{0}_{\mathrm{Total}}-\sigma^{0}_{\mathrm{Total}}\,\nabla\cdot\mathbf{v}$ (148)
$\quad+\nabla\cdot\left(\frac{\bar{\kappa}}{T}\cdot\nabla T\right)+\frac{1}{T^{2}}\nabla T\cdot\bar{\kappa}\cdot\nabla T+\frac{1}{T}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla\mu^{0}_{\Gamma}\cdot\bar{D}\cdot\nabla\mu^{0}_{\Gamma}\,.$
@@
$=\int_{\Omega}\frac{1}{T}\left[\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla T\cdot\bar{\kappa}\cdot\nabla T+\nabla\mu^{a}_{\Gamma}\cdot\bar{D}\cdot\nabla\mu^{a}_{\Gamma}\right]\geq 0\,.$ (149)
@@ 3.4 Metriplectic 2-bracket for the Cahn-Hilliard-Navier-Stokes system
$(F,{G})^{a}_{{H^{a}}}=(F,{H^{a}};{G},{H^{a}})^{a}$ $=\int_{\Omega}\frac{1}{T}\Big{[}\left[T\nabla F_{\mathbf{m}}-F_{{\sigma}^{a}}\nabla\mathbf{v}\right]:\bar{\bar{\Lambda}}:\left[T\nabla G_{\mathbf{m}}-G_{{\sigma}^{a}}\nabla\mathbf{v}\right]$
$+\ \frac{1}{T}\,\big{[}T\nabla F_{{{\sigma}^{a}}}-F_{{\sigma}^{a}}\nabla T\big{]}\cdot\bar{\kappa}\cdot\big{[}T\nabla G_{{{\sigma}^{a}}}-G_{{\sigma}^{a}}\nabla T\big{]}$ (150)
@@ 4 Summary and Conclusions
From the examples presented, it is clear that the 4-bracket formalism can be applied to obtain a wide variety of dynamical systems in various fields. In fact it was recently applied to obtain generalized collision operators in kinetic theory (Sato & Morrison, 2024) and a thermodynamically consistent model for radiation hydrodynamics (Tran et al., 2024).
@@ Appendix A Semidefinite curvature
×$\qquad\times\ \frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}\,\frac{\delta F}{\delta\chi^{\gamma}(z^{\prime\prime})}\frac{\delta G}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\ +\mathrm{\ other\ terms}\,.$
Proof A.1.
@@
Hence, we deduce that $K(F,G)>0$.
Finite-dimensional versions of these two lemmas were first reported in Morrison & Updike (2023).

### 29 Feb 2024 v2

@@ 3.3 Metriplectic 4-bracket for the Cahn-Hilliard-Navier-Stokes system
$\psi(\mathbf{x},t)=(\mathbf{m}(\mathbf{x},t),\rho(\mathbf{x},t),\tilde{c}(\mathbf{x},t),\sigma^{a}(\mathbf{x},t))$ (134)
@@
${\Sigma}(dF,dG)$ $=\nabla F_{{\mathbf{m}}}:\bar{\bar{\Lambda}}_{1}:\nabla G_{{\mathbf{m}}}+\nabla F_{{{\sigma}^{a}}}\cdot\bar{\Lambda}_{2}\cdot\nabla G_{{{\sigma}^{a}}}$
$\hskip 42.67912pt+\nabla\mathcal{L}^{a}_{\tilde{c}}(F)\cdot\bar{\Lambda}_{3}\cdot\mathcal{L}^{a}_{\tilde{c}}(F)\,,$ (136)
@@
Here the tensors $\bar{\bar{\Lambda}}_{1}$ , $\bar{\Lambda}_{2}$ and $\bar{\Lambda}_{3}$ are defined by (59). Then, the 4-bracket reads
$(F,K;G,N)^{a}$ $=$ (138)
$\int_{\Omega}\!\frac{1}{T}\Big{[}\left[K_{{\sigma}^{a}}\nabla F_{\mathbf{m}}-F_{{\sigma}^{a}}\nabla K_{\mathbf{m}}\right]\!\colon\!\!\bar{\bar{\Lambda}}\colon\!\!\left[N_{{\sigma}^{a}}\nabla G_{\mathbf{m}}-G_{{\sigma}^{a}}\nabla N_{\mathbf{m}}\right]$
@@
$+\!\big{[}K_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(F)-F_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(K)\big{]}\!\cdot\!\bar{D}\!\cdot\!\big{[}N_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(G)-G_{{\sigma}^{a}}\mathcal{L}^{a}_{\tilde{c}}(N)\big{]}.$
Observe, with the exception of the last line this bracket is identical to that of (60).
@@
$=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot\big{[}p\mathbf{I}+\lambda_{f}\rho\Gamma{\bm{\xi}}\otimes\nabla c\big{]}$
$\hskip 85.35826pt+\frac{1}{\rho}\nabla\cdot(\bar{\bar{\Lambda}}:\nabla\mathbf{v})\,,$ (140)
$\partial_{t}\rho$ =$=\{\rho,H^{1}\}^{1}+(\rho,H^{1};S^{1},H^{1})^{1}$
$=-\mathbf{v}\cdot\nabla\rho-\rho\,\nabla\cdot\mathbf{v}\,,$ (141)
$\partial_{t}\tilde{c}$ =$=\{\tilde{c},H^{1}\}^{1}+(\tilde{c},H^{1};S^{1},H^{1})^{1}$
$=-\mathbf{v}\cdot\nabla\tilde{c}-\tilde{c}\,\nabla\cdot\mathbf{v}+\nabla\cdot(\bar{D}\cdot\nabla\mu^{1}_{\Gamma})\,,$ (142)
$\partial_{t}\sigma^{1}_{\mathrm{Total}}$ =$=\{\sigma^{1}_{\mathrm{Total}},H^{1}\}^{1}+(\sigma^{1}_{\mathrm{Total}},H^{1};S^{1},H^{1})^{1}$
$=-\mathbf{v}\cdot\nabla\sigma^{1}_{\mathrm{Total}}-\sigma^{1}_{\mathrm{Total}}\,\nabla\cdot\mathbf{v}$
$\quad+\nabla\cdot\left(\frac{\bar{\kappa}}{T}\cdot\nabla T\right)+\frac{1}{T^{2}}\nabla T\cdot\bar{\kappa}\cdot\nabla T$ (143)
$\hskip 28.45274pt+\frac{1}{T}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla\mu^{1}_{\Gamma}\cdot\bar{D}\cdot\nabla\mu^{1}_{\Gamma}\,.$
@@
$=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot\Big{[}\left(p-\lambda_{f}\Gamma^{2}/2\right)\mathbf{I}$
$\hskip 28.45274pt+\lambda_{f}\Gamma{\bm{\xi}}\otimes\nabla c\Big{]}+\frac{1}{\rho}\nabla\cdot(\bar{\bar{\Lambda}}:\nabla\mathbf{v})\,,$ (144)
$\partial_{t}\rho$ =$=\{\rho,H^{0}\}^{0}+(\rho,H^{0};S^{0},H^{0})^{0}$
$=-\mathbf{v}\cdot\nabla\rho-\rho\,\nabla\cdot\mathbf{v}\,,$ (145)
$\partial_{t}\tilde{c}$ =$=\{\tilde{c},H^{0}\}^{0}+(\tilde{c},H^{0};S^{0},H^{0})^{0}$
$=-\mathbf{v}\cdot\nabla\tilde{c}-\tilde{c}\,\nabla\cdot\mathbf{v}+\nabla\cdot(\bar{D}\cdot\nabla\mu^{0}_{\Gamma})\,,$ (146)
$\partial_{t}\sigma^{0}_{\mathrm{Total}}$ =$=\{\sigma^{0}_{\mathrm{Total}},H^{0}\}^{0}+(\sigma^{0}_{\mathrm{Total}},H^{0};S^{0},H^{0})^{0}$
$=-\mathbf{v}\cdot\nabla\sigma^{0}_{\mathrm{Total}}-\sigma^{0}_{\mathrm{Total}}\,\nabla\cdot\mathbf{v}$ (147)
$\quad+\nabla\cdot\left(\frac{\bar{\kappa}}{T}\cdot\nabla T\right)+\frac{1}{T^{2}}\nabla T\cdot\bar{\kappa}\cdot\nabla T$
$\hskip 28.45274pt+\frac{1}{T}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla\mu^{0}_{\Gamma}\cdot\bar{D}\cdot\nabla\mu^{0}_{\Gamma}\,.$
@@
$=\int_{\Omega}\frac{1}{T}\bigg{[}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla T\cdot\bar{\kappa}\cdot\nabla T$
$\hskip 88.2037pt+\nabla\mu^{a}_{\Gamma}\cdot\bar{D}\cdot\nabla\mu^{a}_{\Gamma}\bigg{]}\geq 0\,.$ (148)
@@ 3.4 Metriplectic 2-bracket for the Cahn-Hilliard-Navier-Stokes system
$(F,{G})^{a}_{{H^{a}}}$ $=(F,{H^{a}};{G},{H^{a}})^{a}$ (149)
$=\int_{\Omega}\frac{1}{T}\Big{[}\left[T\nabla F_{\mathbf{m}}-F_{{\sigma}^{a}}\nabla\mathbf{v}\right]:\bar{\bar{\Lambda}}:\left[T\nabla G_{\mathbf{m}}-G_{{\sigma}^{a}}\nabla\mathbf{v}\right]$
$+\ \frac{1}{T}\,\big{[}T\nabla F_{{{\sigma}^{a}}}-F_{{\sigma}^{a}}\nabla T\big{]}\cdot\bar{\kappa}\cdot\big{[}T\nabla G_{{{\sigma}^{a}}}-G_{{\sigma}^{a}}\nabla T\big{]}$
@@ 4 Summary and Conclusions
From the examples presented, it is clear that the 4-bracket formalism can be applied to obtain a wide variety of dynamical systems in various fields. In fact it was recently applied to obtain generalized collision operators in kinetic theory (Sato and Morrison, 2024) and a thermodynamically consistent model for radiation hydrodynamics (Tran et al., 2024).
@@ Appendix A Semidefinite curvature
×$\times\ \frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}\,\frac{\delta F}{\delta\chi^{\gamma}(z^{\prime\prime})}\frac{\delta G}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\ +\mathrm{\ other\ terms}\,.$
Proof 1
@@
Hence, we deduce that $K(F,G)>0$
Finite-dimensional versions of these two lemmas were first reported in Morrison and Updike (2023).